OpenStudy (anonymous):

find the values of the remaining trigonometric funtions a theta from the given information. a) sin theta = 5/13 , cos theta = -12/13 b) cos theta = -3/5 , tan theta < 0

hartnn (hartnn):

good do you want to use standard relations or the triangle method ?

hartnn (hartnn):

you know standard relations like sec x = 1/cos x , tan x = sin x/ cos x and all ? or are you comfortable with drawing a right triangle ?

hartnn (hartnn):

you still there ?

OpenStudy (anonymous):

im still here!

OpenStudy (anonymous):

yea drawing the right triangle

OpenStudy (anonymous):

i want to use the triangle method

hartnn (hartnn):

good, go ahead and draw the triangle

OpenStudy (anonymous):

i need help on how to find the values

hartnn (hartnn):

sin theta = 5/13 you know that sin of angle = opposite side/ hypotenuse right ??? so draw a right triangle using this info., can u ?

OpenStudy (anonymous):



OpenStudy (anonymous):

is that correct?

hartnn (hartnn):

-5 ? side length cannot be negative and which angle is theta ?

OpenStudy (anonymous):

i meant 5 sorry the angle which is theta is sin ? opposite of hypotenuse

hartnn (hartnn):

 got that ^ ??? can you find the missing side ?